# A systematic search for big dippers in ASAS-SN

### v1

@@ Abstract
Variable stars offer a valuable opportunity to study the structure and evolution of stars and stellar populations. For example, intrinsic variables, like RR Lyrae and Cepheid pulsators, are used to study the internal structure of the stars (Marconi et al., 2015; Caputo et al., 2000). Extrinsic variables, like rotational variables and eclipsing binaries, are stars that are variable due to occultaions from other objects or features on the star. Eclipsing binaries are important tools for measuring the masses and radii of stars (Andersen, 1991; Torres et al., 2010; Rowan et al., 2025), and rotational variables can inform models of chromospheric activity and magnetic fields (Cheng et al., 1986), as well as can be used in gyrochronology to determine the age of stellar systems (Barnes, 2007).
@@
Transient surveys generally look for rapidly brightening sources such as supernovae, and variability surveys look for continuously variable sources such as eclipsing binaries, pulsators, and rotational variables. There are few systematic searches for "occasionally" dimming sources. Young stellar objects (YSOs) like T Tauri stars (Bredall et al., 2020) and R Coronae Borealis stars (R Cor Bor stars, Crawford et al., 2025) can show dramatic fading events; however, these events are recurrent. One class of occasionally dimming sources show long, deep, dimming events caused by disks. These objects are often discussed in the context of Aurigae (Carroll et al., 1991) and KH 15D (Kearns & Herbst, 1998). Aurigae has a period of 27.1 years, with dips that reach a depth of mag and last for about 2 years due to an occultaion by a companion B star with a disk (Stefanik et al., 2010). KH 15D has a period of 48 days, but has variable eclipses because of the inclination of its circumbinary disk (Winn et al., 2006; Poon et al., 2021). However, this classification seems to be insufficient to explain the wealth of configurations that can lead to such long and deep events (see Forés-Toribio et al., 2025). OGLE-LMC-ECL-11893 is another object that features eclipses caused by a disk; however, the eclipses have the same shape over many periods, which is not typical for disk eclipsing systems (Dong et al., 2014). Boyajian’s star (Boyajian et al., 2016; Simon et al., 2018) and Zachy’s star (Way et al., 2019) are more examples of objects that have long and deep dimming events. The Via Lactea (VVV) Survey (Minniti et al., 2010) carried out a search for objects with large amplitude dimming and continuous variability in Lucas et al. (2024). Their search discovered 40 YSOs and 21 dipping giant star candidates. These objects are further categorized and characterized in Guo et al. (2024).
@@
The All-Sky Automated Survey for SuperNovae (ASAS-SN, Shappee et al., 2014; Kochanek et al., 2017; Hart et al., 2023) is a transient survey that can also be used to search for long period EBs and dipper stars. It has a baseline of roughly 10 years and, in its current configuration of 20 telescopes on 5 mounts across 4 sites (Chile, Hawaii, South Africa, and Texas), and it achieves a sub-daily cadence across the visible sky in good conditions. ASAS-SN has discovered several "occasional" dimming events. Examples include ASASSN-21co, an eclipsing binary with a period of 11.9 years (Rowan et al., 2021), and ASASSN-23ht, an object with one deep dimming event in the ASAS-SN photometry (Miller et al., 2023). A second class, which we refer to as "big dipper stars", show deep ( mag) dimming events that can last tens to hundreds of days. These objects are distinct from the dippers defined by Ansdell et al. (2016). In the context of this work, dippers are objects with non-recurrent, non-symmetric dimming events that can be long-lasting and have a wide variety of features. These events are likely caused by occultations by extended dusty structures that are produced in a variety of different mechanisms. Examples include ASASSN-24fw, an object with a deep dimming event caused by a cloud of dust (Forés-Toribio et al., 2025; Zakamska et al., 2025), and ASASSN-21qj, a star with a deep, long-lasting event that is thought to be caused by dust formed in a collision of objects orbiting the star (Kenworthy et al., 2023; Marshall et al., 2023).
@@ 2. Methods
We obtained the ASAS-SN light curves of the 5.1 million sources with mag using SkyPatrol v2 (Hart et al., 2023). We use the -band light curves to search for candidates, but include the -band photometry to characterize some of the detected candidates. We improve the calibrations between the ASAS-SN cameras using a damped random walk Gaussian process for interpolation to optimize the camera offsets (see, e.g., Kozłowski et al., 2010). We then designed a pipeline to find all objects that had been serendipitously discovered by ASAS-SN. We searched the light curves for variations with mag. A detected dip was defined as at least one point that satisfied the depth condition. We then removed all stars with light curve dispersions mag. This removes many common variables with high amplitude, (quasi-)continuous variability (e.g., shorter period EBs, YSOs, R Cor Bor stars). This left us with candidates that we visually inspected to identify the strongest candidates and remove false positives. We drop candidates if the dip is only seen in one ASAS-SN camera.
@@
Figure 2 shows examples of false positives rejected during visual inspection. Systematic false positives can be seen in panels (a) through (d) and astrophysical false positives can be seen in panels (e) through (h). Some show continuous variability, but not ”occasional” events. Figure 2a shows the light curve of a star next to a nearby bright star. In Figure 2b there is a group of low quality measurements from a specific camera that triggered its selection. This is the problem leading to the requirement of detections in multiple cameras. Figure 2c shows an example of a failure mode that only occurs near the poles. A source almost always appears in only one field for a given camera, so the light curve intercalibration procedure assumes a single offset for each camera. However, very close to the poles, field rotations can allow a source to appear in two fields for the same camera, leading to problems if the fields need different offsets. Panels (d) through (g) of Figure 2 are variable stars. Panel (d) is an R Coronae Borealis variable and panels (e) through (g) are more common contaminates (contact binaries, semi-regular variables, and YSOs, respectively).
@@
In total, we found 19 new candidates and recover 11 of the 12 previously known systems. J114712621037 was not recovered because the dipping event is shallower than mag and had a very long time scale (1000 days). While we could lower the required magnitude change threshold of , this would have significantly increased the number of targets that would need to be visually inspected. Both the previously known objects and the newly discovered objects are shown on a Gaia color-magnitude diagram (CMD) in Figure 2 (Gaia Collaboration et al., 2016, 2023a) and Table 1 lists the sources, labeled by whether they were discovered earlier as a part of daily ASAS-SN transient search or using our pipeline, along with the Gaia absolute magnitude and colors. We compute extinctions using the mwdust 3-dimensional dust map (Bovy et al., 2016, which is based on, Marshall et al. 2006, Green et al. 2015, and Drimmel et al. 2003), and use distances from Bailer-Jones et al. (2021). Table 2 lists the number of dips found, the maximum dip duration, the epoch of the most recent dip and the maximum dip depth in the -band. The two tables divide the sources into three categories: probable EBs with only a single eclipse, probable EBs with multiple eclipses, and dipper stars.
@@
We checked the -band ASAS-SN data for each of our targets to look for additional events that were missed by our search, as we only applied our automated search to the -band ASAS-SN data. For each of the candidates, we checked the ALLWISE/NEOWISE (Wright et al., 2010; Mainzer et al., 2011) light curves, extracted using tools from Hwang & Zakamska (2020). One source, J183210173432, lacked WISE data due to the proximity of a bright star. We searched VizieR (Ochsenbein et al., 2000) for archival data on each source, in particular for reports of X-ray detections, infrared excesses, and prior variability. We examined LAMOST (Luo et al., 2015) and APOGEE (Eisenstein et al., 2011) spectra when available to search for Balmer (or other) emission lines. We used WISE, 2MASS (Cutri et al., 2003), GALEX (Bianchi et al., 2011), and Gaia (Gaia Collaboration et al., 2023b) photometry to examine the source spectral energy distributions (SEDs), in particular to look for mid-IR emission excesses indicating the presence of circumstellar dust. We visually inspect the SEDs and compare the photometry to stellar models using Castelli & Kurucz (2003) model atmospheres from pystellibs111https://github.com/mfouesneau/pystellibs to identify targets with IR excesses. Finally, for targets flagged as variable stars in other catalogs, we use a Lomb-Scargle periodogram (Lomb, 1976; Scargle, 1982) to determine the period. We also looked for indications of binarity using Gaia.
@@
We used two Gaia statistics to search for evidence of binarity. The Gaia parameter RUWE (renormalized unit weight error) is related to the astrometric goodness of fit and a value higher than 1.4 is a good indicator of binarity (Gaia Collaboration et al., 2016, 2023a). The Gaia parameter rv_amplitude_robust (), is the peak-to-peak amplitude of the Gaia RV measurements after outlier rejection. It can be used to identify binaries without orbital solutions, with systems having km/s very likely to be binaries (Gaia Collaboration et al., 2016, 2023a).
@@ 3. Discussion of each target
where is the semimajor axis and () is the total system mass (in solar masses) and () is the period (in days). We would observe , where is the inclination, but for wide eclipsing binaries with , we can estimate a minimum system mass by comparing and . Assuming the RVs are measured for the more massive star (i.e., the most luminous), can be reduced by increasing the mass ratio and ellipticity can increase it by at most . The amplitude can also be underestimated because the RV measurements do not fully sample orbits longer than the roughly 1000 day (2.8 years) span of the data in Gaia DR3. This also means that once km/s, the Gaia data are no longer sampling the full orbit and Eqn. 1 will begin to increasingly overestimate periods.
@@ 3.1 Single-Dip Eclipsing Binaries
The single-dip eclipsing binaries are objects that have a single symmetric dip in their light curves. The symmetric shape of the dip implies that the occulting object has a regular shape, like a star, rather than an asymmetric, non-uniform, dusty occulter. As only one dip is seen in the light curve, any other eclipses either occur in observational gaps, or the period of the binary is longer than the operational time of ASAS-SN ( years). The deepest of these single eclipses may be dippers, rather than EBs, but they are symmetric and within the limits of EB discoveries. Deeper dips and longer periods become increasingly unlikely to be EBs.
J073234200049 (ASASSN-23ht, Figure 3.1, Miller et al., 2023) is an upper main sequence star with a single dimming event that began on 09/29/2023 and lasted for about 114 days. It reached a depth of mag. This eclipse depth and duration combination is possible for an EB, and the eclipse is close to being symmetric, but it could plausibly be a dipper star instead of an EB.
J175602013135 (ASASSN-24cf, ASAS-SN Transient Page222https://www.astronomy.ohio-state.edu/asassn/transients.html) is on the giant branch and has a single dimming event that began on 22/02/2024 and lasted for about 95 days with a maximum depth of mag. This object is identified as a RS Canum Venaticorum (RS CVn) variable in Gaia (Gaia Collaboration et al., 2016, 2023a). It is difficult for EBs to have this combination of depth and duration, but not impossible.
J223332565552 (ASASSN-23ik, ASAS-SN Transient Page22footnotemark: 2) is a lower main sequence star with a single dimming event that began 15/11/2023 and lasted for about 20 days with a maximum depth of mag. It has been classified as a Persei-type (Algol) eclipsing system, however, there is no reported period (Watson et al., 2006). With an km/s, we can estimate a period of days, using Eqn. 1 and assuming an edge-on circular orbit. However, we see no other events in either the or band light curves.
J073924272916 had a single, mag dimming event that began on 25/02/2023 and lasted for about 27 days. It is a red giant and the event is similar to the event seen in J181752580749 below. ATLAS reports this target as a ”dubious” binary with days (Heinze et al., 2018). With an km/s, the period estimate using Eqn. 1 is days.
@@
J162209444247 had a single, deep mag dimming event on 25/04/2023 that lasted for about 3 days. It is located in the red clump. The ASAS-SN light curve also shows evidence for variability on long timescales, similar to the spotted giants in Petz & Kochanek (2025). After masking the dimming event, we used a Lomb-Scargle periodogram to identify an 11.4 day period, presumably corresponding to the rotation period of the giant. J162209444247 has also been identified as a coronal line emitter in eROSITA (Merloni et al., 2024), with an X-ray luminosity of erg s-1. There is no Gaia orbital solution, but the large RV amplitude ( km/s) suggests that the system is most likely a long-period, chromospherically active, eclipsing binary. The large also implies days (Eqn. 1), which means that the observed Gaia RV variability does not correspond to the long-period binary orbit.
@@
J183606314826 had a single, mag dimming event on 10/03/2022 that lasted for about 11 days. It is located on the giant branch and is similar to J181752580749. AAVSO and ASAS-SN report J183606314826 as a semi-regular variable with days (Henden et al., 2015; Christy et al., 2023). The amplitude of the periodic variability at days is smaller than that of the dip. There is a group of bad points from one camera, creating a cloud below the baseline light curve shortly after the event. This is likely due to an intercalibration issue between the fields for that camera. As discussed above, the km/s implies an orbit long compared to the day span of the Gaia data.
@@
J212132480140 had a single, mag dimming event on 09/09/2021 that lasted for about 58 days. It is located on the giant branch and is similar to J181752580749. Gaia Collaboration et al. (2016, 2023a) reports the detection of short-period variability ( days) for this star, but we do not recover this signal in the ASAS-SN data. The RV amplitude of km/s suggests a long orbit compared to the day span of the Gaia data.
J225702562312 had a single, mag dimming event on 08/06/2019 that lasted for about 6 days. It is located in the red clump and is similar to J175602013135. Gaia reports J225702562312 as a Persei (Algol) eclipsing binary with days (Gaia Collaboration et al., 2016, 2023a). The low km/s implies a long orbit compared to the day span of Gaia.

### v2

@@ Abstract
Variable stars offer a valuable opportunity to study the structure and evolution of stars and stellar populations. For example, intrinsic variables, like RR Lyrae and Cepheid pulsators, are used to study the internal structure of the stars (Marconi et al., 2015; Caputo et al., 2000). Extrinsic variables, like rotational variables and eclipsing binaries, are stars that are variable due to occultations from other objects or features on the star. Eclipsing binaries are important tools for measuring the masses and radii of stars (Andersen, 1991; Torres et al., 2010; Rowan et al., 2025), and rotational variables can inform models of chromospheric activity and magnetic fields (Cheng et al., 1986), and can be used in gyrochronology to determine the age of stellar systems (Barnes, 2007).
Transient surveys generally look for rapidly brightening sources such as supernovae, and variability surveys look for continuously variable sources such as eclipsing binaries, pulsators, and rotational variables. There are few systematic searches for "occasionally" dimming sources. Young stellar objects (YSOs) like T Tauri stars (Bredall et al., 2020) and R Coronae Borealis stars (R Cor Bor stars, Crawford et al., 2025) can show dramatic fading events; however, these events are recurrent.
@@
One class of occasionally dimming sources show long, deep, dimming events caused by disks. These objects are often discussed in the context of Aurigae (Carroll et al., 1991) and KH 15D (Kearns and Herbst, 1998). Aurigae has a period of 27.1 years, with dips that reach a depth of mag and last for about 2 years due to an occultaion by a companion B star with a disk (Stefanik et al., 2010). KH 15D has a period of 48 days, but has variable eclipses because of the inclination of its circumbinary disk (Winn et al., 2006; Poon et al., 2021). However, this classification seems to be insufficient to explain the wealth of configurations that can lead to such long and deep events (see Forés-Toribio et al., 2025). OGLE-LMC-ECL-11893 is another object that features eclipses caused by a disk; however, the eclipses have the same shape over many periods, which is not typical for disk eclipsing systems (Dong et al., 2014). Boyajian’s star (Boyajian et al., 2016; Simon et al., 2018) and Zachy’s star (Way et al., 2019) are more examples of objects that have long and deep dimming events. The Via Lactea (VVV) Survey (Minniti et al., 2010) carried out a search for objects with large amplitude dimming and continuous variability in Lucas et al. (2024). Their search discovered 40 YSOs and 21 dipping giant star candidates. These objects are further categorized and characterized in Guo et al. (2024).
@@
The All-Sky Automated Survey for SuperNovae (ASAS-SN, Shappee et al., 2014; Kochanek et al., 2017; Hart et al., 2023) is a transient survey that can also be used to search for long period EBs and dipper stars. It has a baseline of roughly 10 years and, in its current configuration of 20 telescopes on 5 mounts across 4 sites (Chile, Hawaii, South Africa, and Texas), it achieves a sub-daily cadence across the visible sky in good conditions. ASAS-SN has discovered several "occasional" dimming events. Examples include ASASSN-21co, an eclipsing binary with a period of 11.9 years (Rowan et al., 2021), and ASASSN-23ht, an object with one deep dimming event in the ASAS-SN photometry (Miller et al., 2023). A second class, which we refer to as "big dipper stars", show deep ( mag) dimming events that can last tens to hundreds of days. These objects are distinct from the dippers defined by Ansdell et al. (2016). In the context of this work, dippers are objects with non-recurrent, non-symmetric dimming events that can be long-lasting and have a wide variety of features. These events are likely caused by occultations by extended dusty structures that are produced in a variety of different mechanisms. Examples include ASASSN-24fw, an object with a deep dimming event caused by a cloud of dust (Forés-Toribio et al., 2025; Zakamska et al., 2025), and ASASSN-21qj, a star with a deep, long-lasting event that is thought to be caused by dust formed in a collision of objects orbiting the star (Kenworthy et al., 2023; Marshall et al., 2023).
@@ 2. Methods
We obtained the ASAS-SN light curves of the 5.1 million sources with mag using SkyPatrol v2 (Hart et al., 2023). This magnitude range was selected because it provided a significant sample of unsaturated, fairly high signal-to-noise ratio light curves for an initial search. We use the -band light curves to search for candidates, but include the -band photometry to characterize some of the candidates. We improve the calibrations between the ASAS-SN cameras using a damped random walk Gaussian process for interpolation to optimize the camera offsets (see, e.g., Kozłowski et al., 2010). We then designed a pipeline to find all objects that had been serendipitously discovered by ASAS-SN. We searched the light curves for variations with mag. A detected dip was defined as at least one point that satisfied the depth condition. We then removed all stars with light curve dispersions mag. This removes many common variables with high amplitude, (quasi-)continuous variability (e.g., shorter period EBs, YSOs, R Cor Bor stars). We initially allowed events to be in only one camera. As we started visual inspection, it was clear that we could reduce the false positive rate by requiring dips in multiple cameras with temporal separations of less than one day. With this change, we had candidates, of which were known variables. These were all visually inspected.
@@
Figure 2 shows examples of false positives rejected during visual inspection. Systematic false positives can be seen in panels (a) through (c) and astrophysical false positives can be seen in panels (d) through (g). Some show continuous variability, but not ”occasional” events. Figure 2a shows the light curve of a star too close to a bright star. In many cases this is due to the bleed traits of quite distant saturated stars. These make up of our false positives. In Figure 2b there is a group of problematic data from a specific camera that triggered its selection. This is the problem leading to the requirement of detections in multiple cameras. Roughly of our false positives are due to this issue. Figure 2c shows an example of a failure mode that only occurs near the poles. A source almost always appears in only one field for a given camera, so the light curve intercalibration procedure assumed a single offset for each camera. However, close to the equatorial poles, field rotations can allow a source to appear in two fields for the same camera, leading to problems if the fields need different offsets. This error caused of our false positives. Panels (d) through (g) of Figure 2 are variable stars. Panel (d) is an R Coronae Borealis variable and panels (e) through (g) are more common contaminates (contact binaries, semi-regular variables, and YSOs, respectively). Known variables made up of the false positives.
@@
In total, we found 19 new candidates and recover 11 of the 12 previously known systems. J114712621037 was not recovered because its dipping event is shallower than mag and had a very long time scale (1000 days). While we could lower the required magnitude change threshold of , this would have significantly increased the number of targets that would need to be visually inspected. Both the previously known objects and the newly discovered objects are shown on a Gaia color-magnitude diagram (CMD) in Figure 2 (Gaia Collaboration et al., 2016, 2023b) and Table 1 lists the sources, labeled by whether they were discovered earlier as a part of daily ASAS-SN transient search or using our pipeline, along with the Gaia absolute magnitude and colors. We compute extinctions using the mwdust 3-dimensional dust map (Bovy et al., 2016, which is based on, Marshall et al. (2006), Green et al. (2015), and Drimmel et al. (2003)), and use distances from Bailer-Jones et al. (2021). Table 2 lists the number of dips found, the maximum dip duration, the epoch of the most recent dip and the maximum dip depth in the -band. The two tables divide the sources into three categories: probable EBs with only a single eclipse, probable EBs with multiple eclipses, and dipper stars. The light curves for each of the final candidates that had passed the visual inspection can be found in a supplemental folder accompanying this work. This folder includes zoomed in sections of each light curve corresponding to where the pipeline was triggered.
@@
We checked the earlier -band ASAS-SN data for each of our final candidates to look for additional dimming events. We also checked their ALLWISE/NEOWISE (Wright et al., 2010; Mainzer et al., 2011) light curves, extracted using tools from Hwang and Zakamska (2020). One source, J183210173432, lacked WISE data due to the proximity of a bright star. We searched VizieR (Ochsenbein et al., 2000) for archival data on each source, in particular for reports of X-ray detections, infrared excesses, and prior variability. We examined LAMOST (Luo et al., 2015) and APOGEE (Eisenstein et al., 2011) spectra when available to search for Balmer (or other) emission lines. We used WISE, 2MASS (Cutri et al., 2003), GALEX (Bianchi et al., 2011), and Gaia (Gaia Collaboration et al., 2023a) photometry to examine the source spectral energy distributions (SEDs), in particular to look for mid-IR emission excesses relative to Castelli and Kurucz (2003) model stellar atmospheres from pystellibs111https://github.com/mfouesneau/pystellibs indicating the presence of circumstellar dust. Finally, for targets flagged as variable stars in other catalogs, we use a Lomb-Scargle periodogram (Lomb, 1976; Scargle, 1982) to determine the period.
@@
We used two Gaia statistics to search for evidence of binarity. The Gaia parameter RUWE (renormalized unit weight error) is related to the astrometric goodness of fit and a value higher than 1.4 is a good indicator of binarity (Gaia Collaboration et al., 2016, 2023b). The Gaia parameter rv_amplitude_robust (), is the peak-to-peak amplitude of the Gaia RV measurements after outlier rejection. It can be used to identify binaries without orbital solutions, with systems having km/s very likely to be binaries (Gaia Collaboration et al., 2016, 2023b).
@@ 3. Discussion of each target
where is the semimajor axis and () is the total system mass (in solar masses) and () is the period (in days). We observe , where is the inclination, but for wide eclipsing binaries with , we can estimate a minimum system mass by comparing and . Assuming the RVs are measured for the more massive star (i.e., the most luminous), can be reduced by increasing the mass ratio and ellipticity can increase it by at most . The amplitude can also be underestimated because the RV measurements do not fully sample orbits longer than the roughly 1000 day (2.8 years) span of the data in Gaia DR3. Once the Gaia data are no longer sampling the full orbit, Eqn. 1 will begin to increasingly overestimate periods. Another issue is that once gets close to km/s, much of it could be created by noise in the RV measurements, leading to Eqn. 1 to underestimate the period.
@@ 3.1 Single-Dip Eclipsing Binaries
The single-dip eclipsing binaries are objects that have a single symmetric dip in their light curves. The shapes of the multi-camera eclipse light curve, such as having a symmetric ingress and egress, are used to visually confirm an eclipse. The symmetric shape of the dip implies that the occulting object has a regular shape, like a star, rather than an asymmetric or non-uniform dusty occulter. As only one dip is seen in the light curve, any other eclipses either occur in observational gaps, or the period of the binary is longer than the operational time of ASAS-SN ( years). The deepest of these single eclipses may be dippers, rather than EBs, but they are symmetric and have depths within the limits possible for EBs (see Forés-Toribio et al., 2025). Deeper dips and longer periods become increasingly unlikely to be EBs. We folded the light curves as a function of period to find the minimum period where prior eclipses could all fit in seasonal gaps or prior to the start of the survey. They are presented in Table 2.
J073234200049 (ASASSN-23ht, Figure 3.1, Miller et al., 2023) is an upper main sequence star with a single dimming event that began on 09/29/2023 and lasted for about 114 days. It reached a depth of mag. This eclipse depth and duration combination are possible for an EB, and the eclipse is close to being symmetric, but it could plausibly be a dipper star instead of an EB. The minimum period based on the light curve is years.
J175602013135 (ASASSN-24cf, ASAS-SN Transient Page222https://www.astronomy.ohio-state.edu/asassn/transients.html) is on the giant branch and has a single dimming event that began on 22/02/2024 and lasted for about 95 days with a maximum depth of mag. This object is identified as a RS Canum Venaticorum (RS CVn) variable in Gaia (Gaia Collaboration et al., 2016, 2023b). It is difficult for EBs to have this combination of depth and duration, but not impossible.
J223332565552 (ASASSN-23ik, ASAS-SN Transient Page22footnotemark: 2) is a lower main sequence star with a single dimming event that began 15/11/2023 and lasted for about 20 days with a maximum depth of mag. It has been classified as a Persei-type (Algol) eclipsing system, however, there is no reported period (Watson et al., 2006). With an km/s, we can estimate a period of days, using Eqn. 1 and assuming an edge-on circular orbit. However, we see no other events in either the or band light curves with an estimated minimum period of roughly 1700 days. This minimum period is close to two times the estimate using Eqn. 1, however, the is close to the threshold of km/s where underestimates of the period become more likely.
J073924272916 had a single, mag dimming event that began on 25/02/2023 and lasted for about 27 days. It is a red giant and the event is similar to the event seen in J181752580749 below. ATLAS reports this target as a ”dubious” binary with days (Heinze et al., 2018). With an km/s, the period estimate using Eqn. 1 is days. Which is likely an underestimate. From folding the light curve we estimate a minimum period of days.
@@
J162209444247 had a single, deep mag dimming event on 25/04/2023 that lasted for about 3 days. It is located in the red clump. The ASAS-SN light curve also shows evidence for variability on long timescales, similar to the spotted giants in Petz and Kochanek (2025). After masking the dimming event, we used a Lomb-Scargle periodogram to identify an 11.4 day period, presumably corresponding to the rotation period of the giant. J162209444247 has also been identified as a coronal line emitter in eROSITA (Merloni et al., 2024), with an X-ray luminosity of erg s-1. There is no Gaia orbital solution, but the large RV amplitude ( km/s) suggests that the system is most likely a long-period, chromospherically active, eclipsing binary. The large also implies days (Eqn. 1), which means that the observed Gaia RV variability does not correspond to the long-period binary orbit.
@@
J183606314826 had a single, mag dimming event on 10/03/2022 that lasted for about 11 days. It is located on the giant branch and is similar to J181752580749. AAVSO and ASAS-SN report J183606314826 as a semi-regular variable with days (Henden et al., 2015; Christy et al., 2023). The amplitude of the periodic variability at days is smaller than that of the dip. There is a group of bad points from one camera, creating a cloud below the baseline light curve shortly after the event. This is likely due to an intercalibration issue between the fields for that camera. As discussed above, the km/s implies an orbit long compared to the day span of the Gaia data. This is in agreement with the minimum period estimate of days.
@@
J212132480140 had a single, mag dimming event on 09/09/2021 that lasted for about 58 days. It is located on the giant branch and is similar to J181752580749. Gaia Collaboration et al. (2016, 2023b) reports the detection of short-period variability ( days) for this star, but we do not recover this signal in the ASAS-SN data. The RV amplitude of km/s suggests a long orbit compared to the day span of the Gaia data.
J225702562312 had a single, mag dimming event on 08/06/2019 that lasted for about 6 days. It is located in the red clump and is similar to J175602013135. Gaia reports J225702562312 as a Persei (Algol) eclipsing binary with days (Gaia Collaboration et al., 2016, 2023b). The low km/s implies a long orbit compared to the day span of Gaia.